OpenStudy (anonymous):

Write the equation for a circle with its center at (2,-5) and a radius of 6. a.(x + 2)2 + (y - 5)2 = 6 b.(x + 2)2 + (y - 5)2 = 36 c.(x - 2)2 + (y + 5)2 = 6 d.(x - 2)2 + (y + 5)2 = 36

OpenStudy (nenadmatematika):

(x-p)^2+(y-q)^2=R^2 is the equation of the circle where p and q are coordinates of center and R is radius...would you know how to write it now?

OpenStudy (nenadmatematika):

no? (x-2)^2+(y+5)^2=36 so d is your answer

OpenStudy (anonymous):

i think i do . thank you

OpenStudy (nenadmatematika):

you're welcome :D